OpenStudy (anonymous):

How to simplify sqrt((1+sqrt3/2) / 2) ?

OpenStudy (anonymous):

\[\sqrt{(1+\sqrt{3}/2})/2\] I know I need to get rid of the denominators and bring something out of the radical but I'm not sure how to do that

OpenStudy (anonymous):

it is all under the radical?

OpenStudy (misty1212):

\[\huge \sqrt{\frac{1+\frac{\sqrt{3}}{2}}{2}}\]

OpenStudy (anonymous):

multiply top and bottom inside the radical by 2 to clear the compound fraction

OpenStudy (anonymous):

you get \[\sqrt{\frac{2+\sqrt3}{4}}\]

OpenStudy (anonymous):

then take the square root of the denominator and end with \[\frac{\sqrt{2+\sqrt3}}{2}\] that is the most you can do

OpenStudy (anonymous):

okay then thank you